40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Mar 8 21:48:46 2012, Frank, Koji, DailyProgress, Seismic, air spring isolation installation finished IMG068.jpgIMG070.jpgair_spring_vertical_tf.png
    Reply  Sat Mar 10 18:42:42 2012, tara, DailyProgress, BEAT, beat- with air spring isolation  beat_2012_03_10.pngbeat_2012_03_10.png
Message ID: 877     Entry time: Thu Mar 8 21:48:46 2012     Reply to this: 878
Author: Frank, Koji 
Type: DailyProgress 
Category: Seismic 
Subject: air spring isolation installation finished 

got some T-connectors and connected all four springs to the lab air supply. The pressure of 32PSI is barely enough to float the chamber. The side with the horizontal extra flange is almost too heavy to float, but it's OK. The other side is much better. But overall the chamber is pretty good leveled. Tara re-aligned the laser to both cavities and the beam height is ~6.5in as designed.

IMG068.jpg IMG070.jpg

we measured the TF from the table to the vacuum chamber. Sensor is attached to the top of the ion pump in the center (left picture), the other one on the aluminum block next to the chamber (see picture). Resonance frequency is at 5.19Hz and has a transmissibility of ~10. Both numbers are slightly higher than expected from the datasheet.

air_spring_vertical_tf.png

ELOG V3.1.3-